OpenStudy (anonymous):

Solve for x where x is a real number in the equation below. (show work please)

OpenStudy (anonymous):

\[27^{2x}=9^{x-3}\]

OpenStudy (lgbasallote):

Note: \[27 = 3^3\] \[9 = 3^2\] so \[\large 27^{2x} = 9^{x-3}\] \[\large \implies (3^3)^{2x} = (3^2)^{x-3}\] \[\large \implies 3^{6x} = 3^{2(x-3)}\] \[\large \implies 6x = 2(x-3)\] \[\large \implies 6x = 2x - 6\] \[\large \implies 6x - 2x = -6\] \[\large 4x = -6\] \[\large \implies x = -\frac 64\] \[\large \implies x = -\frac 32\] does that help?

OpenStudy (anonymous):

Yes it did. Im just horrible with rules of exponents. Thank you for the help.